OpenStudy (anonymous):

What is the value of the expression 7^–3? 7 to the power of negative 3. plz help

OpenStudy (anonymous):

Help plz!! I'm horrible at math

OpenStudy (anonymous):

If you know the basic rule of algabra ,for example a^-b=1/(a^b) then you can easilly get the answer .So if you apply this rule in this case, you will get 1/(7^3)=1/343=0.00291545

OpenStudy (anonymous):

Thank you sooo much my teacher said this might show up on my TCAP tests